Huang Industries is considering a proposed project whose estimated NPV is $12 million. This estimate assumes that economic conditions will be average. However the CFO realizes that conditions could be better or worse, so she performed a scenario analysis and obtained these results: Economic Scenario Probability of Outcome NPV 0.05 0.20 0.50 0.20 0.05 - $54 million - $18 million 12 mlion 18 milion 30 million Recession Below average Average Above average Boom Calculate the projects expected NPV, standard deviation, and coefficient of variation. Round your answers to two decimal places. Enter your answers for the projects expected NPV and standard deviation in millions. For example, an answer of $13,000,000 should be entered as 13 million E(NPV) million CV
Answer #1
Economic Scenario Probability (P) NPV($ in million) Probability*NPV Deviation form expected NPV (D) PD^2
Recession 0.05 -54 -2.70 -58.80 172.87
Below average 0.20 -18 -3.60 -22.80 103.97
Average 0.50 12 6.00 7.20 25.92
Above average 0.20 18 3.60 13.20 34.85
Boom 0.05 30 1.50 25.20 31.75

Expected NPV = sum(Probability*NPV)

= -2.7-3.6+6+3.6+1.5

= 4.80 million

Variance = Sum(PD^2)

= 172.87+103.97+25.92+34.85+31.75

= 369.36 million

Standard Deviation = Squareroot(Variance)

= 369.36 million^.5

= 19.22 million

Coefficient of Variation = Standard Deviation/ Expected NPV

= 19.22 million / 4.80 million

= 4.00